OpenStudy (anonymous):

A jetliner is traveling east from Seattle to Chicago, a distance of 1,750 miles. The jetliner travels at an average air speed of 620 mph relative to the air. At the jetliner's altitude, the wind is blowing 46 mph to the east. How long does the journey take? a] 2.82 hours b]0.354 hours c] 3.05 hours d]2.63 hours

OpenStudy (anonymous):

I think it should be 2.63 hours....

OpenStudy (anonymous):

:] thank you

OpenStudy (anonymous):

Vjet - Vair = Vrelative of jet wrt air find Vjet

OpenStudy (anonymous):

Its because, the plane is going at a speed of 620 mph, and wind is flowing at a speed of 46 mph in the same direction.... so vjet=620+46=666mph 1750/666=2.63 hours...